# Refining Concentration for Gaussian Quadratic Chaos

### v2

@@ II Analysis of the results in Subsection I-B
II-A Improved LM Inequality
Consider the improved LM inequality given in (14) in Proposition 2. It is easily checked that $\Lambda(t,a)<0$ for all $t>0$ and $\frac{2}{3}<a\leq 1$ and hence, the bound in (14) is nontrivial. Let us denote the ratio $\frac{\alpha t}{\beta}$ by $\rho$, i.e.,
@@ II-A Improved LM Inequality
As mentioned earlier in Section I, the original LM inequality itself is recovered from (14) by the choices $a=b=1$ and $c=\infty$ and hence, it is given by
@@
Once again, using the fact that $\Lambda(t,a)$ as a function of $a$ is strictly increasing over the interval $[\hat{a}_{opt},1]$, we conclude that the augmented LM bound in (43) is tighter than the classic LM bound in (45). We demonstrate the improvement through an example. Let $n=3$ and the positive-semidefinite matrix $A$ be given by [^12] The positive-semidefinite matrix in (46) is randomly generated.
@@
II-B Beyond HW Inequality
Next, we investigate the sequence of bounds presented in Proposition 4. Recall that the choice $m=1$ recovers the classic HW inequality. Two initial remarks are in order.
@@ II-B Beyond HW Inequality
    As we mentioned earlier, the upper bound $e^{\Lambda_{m}(t,b)}$ must be minimized over $10<b<$ 1. Such minimization results in meaningful bounds in the sense that $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}\leq 1$. This is due to the simple observation that $\lim_{b\to 0^{+}}\Lambda_{m}(t,b)=0$.
@@
    The coefficient $\kappa_{m}(b)$ in the exponent achieves a global maximum value somewhere inside the interval $10<b<$ 1. Since $\lim_{b\to 0^{+}}\theta_{m}(b)=\theta_{m}(0)=\frac{1}{m+1}$ and $\lim_{b\to 1^{-}}\theta_{m}(b)=\infty$, we have $\lim_{b\to 0^{+}}\kappa_{m}(b)=\lim_{b\to 1^{-}}\kappa_{m}(b)=0$. Fig. 3 plots $\kappa_{m}(b)$ in terms of $b$ for several values of $m$.
    Figure 3: Plots of $\kappa_{m}(b)$ in terms of $10<b<$ 1 for $m=1,2,5,10$.
@@
Figure 4: Panel (a) presents plots of the upper bound $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for $t=5$ and $m=1,2,3,4,5$. We see that the infimum $\inf_{0<b<1}e^{\Lambda_{m}(5,b)}$ is the smallest for $m=1$, i.e., HW inequality offers the tightest upper bound. Panel (b) presents similar plots for $t=10$ and $m=1,5,9,13,17$. We see that the achieved infimum value decreases as $m$ grows.
@@
Fig. 4 in panel (a) presents plots of $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for $t=5$ and $m=1,2,3,4,5$. We see that the achieved infimum value $\inf_{0<b<1}e^{\Lambda_{m}(5,b)}$ is the smallest when $m=1$, i.e., HW inequality offers the best bound. In contrast, panel (b) shows plots of $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for the larger tail parameter $t=10$ and $m=1,5,9,13,17$. We see that the achieved infimum value occurs when $m=17$, i.e, when $m$ is the largest. This is an indication of a phase transition in the behaviour of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of the pair of parameters $(m,t)$. As stated in Conjecture 1 in Subsection I-B, there exists a critical threshold for the tail parameter denoted by $t_{c}$ such that if $t<t_{c}$, then $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest when $m=1$ and if $t>t_{c}$, then $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest when $m$ grows to infinity. To better observe this phase transition, Fig. 5 offers plots of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of $1\leq m\leq 20$ for the values $5\leq t\leq 10$ in steps of $0.2$, i.e., $t=5,5.2,5.4,\cdots,9.8,10$. Panel (b) shows the optimum $m$ denoted by $m_{opt}$ over the domain $1\leq m\leq 20$ at which $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest. We have $m_{opt}=1$ for $t\leq 7$ and $m_{opt}=20$ for $t\geq 7.2$. This tells us that if Conjecture 1 is true, then the critical threshold must satisfy $t_{c}\leq 7.2$. In Corollary 3 in below, we will present a theoretical upper bound on $t_{c}$.
We continue by stating three corollaries to Proposition 4. The first one is a loosened version of the upper bound in (21) which no longer involves minimization over the parameter $10<b<$ 1.
@@
Figure 5: Panel (a) presents plots of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of $1\leq m\leq 20$ for $5\leq t\leq 10$ in steps of $0.2$. Panel (b) shows the optimum $m$ denoted by $m_{opt}$ over the domain $1\leq m\leq 20$ at which $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest. If Conjecture 1 is true, then these plots imply that $t_{c}\leq 7.2$.
Corollary 1.
Let $A$ be a symmetric matrix. For every $\epsilon>0$ and integer $m\geq 2$,
$\Pr(\boldsymbol{\Delta}>t)\leq\inf_{0<b<1}e^{\Lambda_{m}(t,b)}\leq(1+\epsilon)\exp\bigg{(}-\tilde{\kappa}_{m}\min\bigg{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg{\}}\bigg{)},$ (48)
where $\tilde{\kappa}_{m}>0$ is a constant that depends only on $m,n$ and $\epsilon$.
Proof.
Consider the upper bound in (21). Choose $10<b<$ 1 such that
$\frac{n}{2}\sum_{k=2}^{m}\frac{b^{k}}{k}=\ln(1+\epsilon).$ (49)
Denote this value of $b$ by $\tilde{b}$ which only depends on $m,n$ and $\epsilon$. Then
$\Lambda_{m}(t,\tilde{b})=\ln(1+\epsilon)-\kappa_{m}(\tilde{b})\min\bigg{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg{\}}$ (50)
and
$\Pr(\boldsymbol{\Delta}>t)\leq e^{\Lambda_{m}(t,\tilde{b})}=(1+\epsilon)\exp\bigg{(}-\kappa_{m}(\tilde{b})\min\bigg{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\bigg{\}}\bigg{)}.$ (51)
This verifies (48) with
$\tilde{\kappa}_{m}=\kappa_{m}(\tilde{b})=\frac{m}{2(m+1)}\min\bigg{\{}\frac{1}{((m+1)\theta_{m}(\tilde{b}))^{\frac{1}{m}}},\tilde{b}\bigg{\}}.$ (52)
The upper bound in (48) is futile for sufficiently small values of $t$ and only becomes a nontrivial bound if $t$ is large enough so that the right side in (48) becomes smaller than one. Moreover, a smaller $\epsilon$ would result in a poorer upper bound for larger values of $t$. This is due to the fact that a smaller $\epsilon$ yields a smaller $\tilde{b}$ as the solution for $b$ in (49) and hence, the coefficient $\tilde{\kappa}_{m}$ in (52) will be smaller. In other words, a larger $\epsilon$ generates a looser bound for smaller values of $t$, however, it will later produce a tighter upper bound when $t$ grows sufficiently large thanks to a larger coefficient $\tilde{\kappa}_{m}$ that sits behind $\min\Big{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\Big{\}}$ in the exponent in (48). Fig. 6 presents plots of the upper bound $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in Proposition 4 for $m=1$ (HW inequality) and $m=20$ in terms of $0\leq t\leq 40$ as well as the upper bounds offered in Corollary 1 for $m=20$ and $\epsilon=0.5,1,2$. We included HW inequality in the plots for comparison.
Figure 6: Plots of the upper bound $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in Proposition 4 for $m=1$ (HW inequality) and $m=20$ in terms of $0\leq t\leq 40$ as well as the weaker upper bounds offered in Corollary 1 for $m=20$ and $\epsilon=0.5,1,2$. A larger $\epsilon$ generates a looser upper bound for smaller values of $t$ and a tighter upper bound when $t$ grows larger. This is due to the fact that a larger $\epsilon$ results in a larger value for the coefficient $\tilde{\kappa}_{m}$ that sits behind $\min\Big{\{}\frac{t^{1+\frac{1}{m}}}{\|A\|_{m+1}^{1+\frac{1}{m}}},\frac{t}{\|A\|}\Big{\}}$ in the exponent in (48).
@@
For our second corollary, we study the behaviour of the bounds in Proposition 4 as $m$ grows to infinity. This will result in what we refer to as the $m_{\infty}$-inequality.
@@
The first inequality in (53) is trivial. For the second inequality, we note that [^14] If $f_{n}:A\to\mathbb{R}$ is an arbitrary sequence of functions on a set $A$ of real numbers, then $\inf_{a\in A}f_{n}(a)\leq f_{n}(x)$ for every $x\in A$ and $n\geq 1$. Taking the limit inferior over $n$ from both sides, $\liminf_{n\to\infty}\inf_{a\in A}f_{n}(a)\leq\liminf_{n\to\infty}f_{n}(x)$. This says $\liminf_{n\to\infty}\inf_{a\in A}f_{n}(a)$ is a lower bound on the set of numbers $\liminf_{n\to\infty}f_{n}(x)$ for $x\in A$. Thus, $\liminf_{n\to\infty}\inf_{a\in A}f_{n}(a)\leq\inf_{x\in A}\liminf_{n\to\infty}f_{n}(x)$ as the infimum of a set is its largest lower bound.
@@
for all $t\geq 0$. The limit $\lim_{m\to\infty}\Lambda_{m}(t,b)$ exists. To see this, note that
@@
where we have used the MacLaurin expansion $-\ln(1-b)=\sum_{k=1}^{\infty}\frac{b^{k}}{k}$ for $10<b<$ 1. By (57), (58) and (59),
@@
Our third and last corollary enables us to find an upper bound on the critical threshold $t_{c}$ in Conjecture 1.
@@
Let $\hat{t}_{c}$ be the unique positive solution for $t$ in the equation
$\Big{(}1+\frac{t}{n\|A\|}\Big{)}^{\frac{n}{2}}e^{-\frac{(\frac{1}{2}-\kappa)t}{\|A\|}}=1,$ (63)
where $\kappa$ is the constant in HW inequality given in Proposition 1. If Conjecture 1 is true, then $t_{c}\leq\hat{t}_{c}$.
@@
$g(t)=\limsup_{m\to\infty}\inf_{0<b<1}e^{\Lambda_{m}(t,b)}.$ (65)
By definition, the critical threshold $t_{c}$ is given by
$t_{c}=\sup\{t>0:g(t)\geq f(t)\}.$ (66)
@@
$t_{c}\leq\hat{t}_{c}=\sup\Big{\{}t>0:\Big{(}1+\frac{t}{n\|A\|}\Big{)}^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}}\geq e^{-\frac{\kappa t}{\|A\|}}\Big{\}}.$ (67)
As an application of Corollary 3, let us consider the symmetric matrix $A$ in (47). Using the plots in Fig. 5, we concluded that $t_{c}\leq 7.2$. Using Corollary 3, we get the sharper bound $t_{c}\leq\hat{t}_{c}=6.191$.
II-C Improvement over HW inequality for both all small $t$ and all large $t$
Here, we look at the twin inequality to HW inequality in Proposition 5. For a numerical study, let us consider the symmetric matrix $A$ in (47). Panels (a) and (b) in Fig. 7 compare the bound in (28) with HW inequality for $0\leq t\leq 2$ and $5\leq t\leq 14$, respectively. We observe that inequality (28) is tighter than HW inequality for both sufficiently small and large values of the tail parameter $t$. As briefly pointed out in Section I, this advantage can be verified by writing asymptotic expansions for the two functions $\eta_{1}(t)$ and $\eta_{2}(t)$ in (26) and (27), respectively. One can easily verify that
@@ II-C Improvement over HW inequality for both all small $t$ and all large $t$
Figure 7: Inequality (28) in Proposition 5 is tighter than HW inequality for both sufficiently small and large values of $t$. For the symmetric matrix $A$ given in (47), Panel (a) and Panel (b) compare the two bounds for $0\leq t\leq 2$ and $5\leq t\leq 14$, respectively. We have also included the $m_{\infty}$-bound in Corollary 2 for comparison.
@@
This shows that $\min\{\eta_{1}(t),\eta_{2}(t)\}$ behaves similarly to the term $\min\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\}$ for sufficiently small and large values of $t$. Since the coefficient $\kappa^{\prime}\approx 0.187$ that sits behind $\min\{\eta_{1}(t),\eta_{2}(t)\}$ is larger than the coefficient $\kappa\approx 0.145$ behind $\min\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\}$, it follows that (28) is tighter than HW inequality for those extreme values of $t$.
@@
Let the matrix $A$ be positive-semidefinite. We present four upper bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through the dimension $n$ and the operator norm $\|A\|$. The first candidate is the $m_{\infty}$-bound in Corollary 2 which in fact holds for every symmetric matrix $A$. For the second and third candidates, we consider relaxed versions of HW inequality and LM inequality, respectively. We know
@@
The exponents in HW inequality and LM inequality are increasing in terms of $\|A\|_{2}$. This is clear in case of HW inequality and its relaxed version in (29) immediately follows. As for LM inequality in (5), note that $\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2}$ is positive and decreasing in terms of $\|A\|_{2}$ and hence, $-(\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2})^{2}$ becomes increasing in terms of $\|A\|_{2}$. This together with (77) gives
@@
Factoring $\sqrt{n}\|A\|$ in $\sqrt{n\|A\|^{2}+2\|A\|t}-\sqrt{n}\|A\|$ results in the relaxed LM inequality in (30). As stated in Proposition 6, the $m_{\infty}$-bound is always tighter than both the relaxed HW and the relaxed LM inequalities. One can also consider the possibility of relaxing the augmented and optimal LM inequalities in (43) and (44), respectively. This requires that we replace the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$ in (36) by $\frac{t\|A\|}{(\sqrt{n}\|A\|)^{2}}=\frac{t}{n\|A\|}$. Numerical results indicate that the $m_{\infty}$-bound is still tighter that these relaxed inequalities and hence, we do not pursue this direction.
@@
[…]
@@
Fig. 8 in panels (a) and (b) presents graphs of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for $n=8$ and $n=10$, respectively. Approximate values for the numbers $r_{n},r^{\prime}_{n}$ are also provided on these plots. Fig. 9 compares all four candidates, namely, the $m_{\infty}$-bound, the relaxed HW bound, the relaxed LM bound and the $\chi^{2}$-bound for $n=8$ and $n=10$. Note that these four bounds are entirely functions of the dimension $n$ and the ratio $r$ in (88).
Figure 9: Comparison between all four candidates, namely, the $m_{\infty}$-bound, the relaxed HW bound, the relaxed LM bound and the $\chi^{2}$-bound for $n=8$ and $n=10$. The formulas for these four bounds are entirely functions of the dimension $n$ and the ratio $r$ in (88).
@@
[…]
This upper bound is too loose to be useful. In fact, it is always weaker than the $m_{\infty}$-bound in (89) for every $n\geq 2$. To see this, call $x=\frac{1+r}{n}\geq 1$. The bound in (94) is looser than that in (89) if and only if
$\iff$ $\Big{(}1+\frac{nx-1}{n}\Big{)}^{\frac{n}{2}}<\frac{1}{\sqrt{e}}(ex)^{\frac{n}{2}}$
$\stackrel{{\scriptstyle(*)}}{{\iff}}$ $1+\frac{nx-1}{n}<\frac{1}{e^{\frac{1}{n}}}ex$
$\iff$ $1-\frac{1}{n}+x<e^{1-\frac{1}{n}}x$
$\iff$ $1-\frac{1}{n}<(e^{1-\frac{1}{n}}-1)x,$
where in ()(*)( ∗ ) we raised both sides to the power of $\frac{2}{n}$. Using the inequality $z<e^{z}-1$ for $z\neq 0$, we have $11-\frac{1}{n}<e^{1-\frac{1}{n}}-$ 1 for every $n\geq 2$. Therefore, (95) is certainly true due to $x\geq 1$.

### v3

@@ II Analysis of the results in Subsection I-B
II-A Improving LMI
Consider the improved LMI given in (17) in Proposition 2. It is easily checked that $\Lambda(t,a)<0$ for all $t>0$ and $\frac{2}{3}<a\leq 1$ and hence, the bound in (17) is nontrivial. Let us denote the ratio $\frac{\alpha t}{\beta}$ by $\rho$, i.e.,
@@ II-A Improving LMI
As mentioned earlier in Section I, the original LMI itself is recovered from (17) by the choices $a=b=1$ and $c=\infty$ and hence, it is given by
@@
Once again, using the fact that $\Lambda(t,a)$ as a function of $a$ is strictly increasing over the interval $\hat{a}_{opt},1]$, we conclude that the augmented LM bound in (47) is tighter than the classic LM bound in (49). We demonstrate the improvement through an example. Let $n=3$ and the positive-semidefinite matrix $A$ be given by [^12] The positive-semidefinite matrix in (50) is randomly generated.
@@
II-B Beyond HWI
Next, we investigate the sequence of bounds presented in Theorem 1. Recall that the choice $m=1$ recovers the classic HWI. Two initial remarks are in order.
@@ II-B Beyond HWI
    As we mentioned earlier, the upper bound $e^{\Lambda_{m}(t,b)}$ must be minimized over $0<b<1$. Such minimization results in meaningful bounds in the sense that $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}\leq 1$. This is due to the simple observation that $\lim_{b\to 0^{+}}\Lambda_{m}(t,b)=0$.
@@
    The coefficient $\kappa_{m}(b)$ in the exponent achieves a global maximum value somewhere inside the interval $0<b<1$. Since $\lim_{b\to 0^{+}}\theta_{m}(b)=\theta_{m}(0)=\frac{1}{m+1}$ and $\lim_{b\to 1^{-}}\theta_{m}(b)=\infty$, we have $\lim_{b\to 0^{+}}\kappa_{m}(b)=\lim_{b\to 1^{-}}\kappa_{m}(b)=0$. Fig. 3 plots $\kappa_{m}(b)$ in terms of $b$ for several values of $m$.
    Figure 3: Plots of $\kappa_{m}(b)$ in terms of $0<b<1$ for $m=1,2,5,10$.
@@
Figure 4: Panel (a) presents plots of the upper bound $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for $t=5$ and $m=1,2,3,4,5$. We see that the infimum $\inf_{0<b<1}e^{\Lambda_{m}(5,b)}$ is the smallest for $m=1$, i.e., HWI offers the tightest upper bound. Panel (b) presents similar plots for $t=10$ and $m=1,5,9,13,17$. We see that the achieved infimum value decreases as $m$ grows.
@@
Fig. 4 in panel (a) presents plots of $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for $t=5$ and $m=1,2,3,4,5$. We see that the achieved infimum value $\inf_{0<b<1}e^{\Lambda_{m}(5,b)}$ is the smallest when $m=1$, i.e., HWI offers the best bound. In contrast, panel (b) shows plots of $e^{\Lambda_{m}(t,b)}$ in terms of $b$ for the larger tail parameter $t=10$ and $m=1,5,9,13,17$. We see that the achieved infimum value occurs when $m=17$, i.e, when $m$ is the largest. This is an indication of a phase transition in the behaviour of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of the pair of parameters $(m,t)$. As stated in Conjecture 1 in Subsection I-B, there exists a critical threshold $\tau_{c}$ such that if $t<\tau_{c}$, then $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest when $m=1$ and if $t>\tau_{c}$, then $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest when $m$ grows to infinity. To better observe this phase transition, Fig. 5 offers plots of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of $1\leq m\leq 20$ for the values $5\leq t\leq 10$ in steps of $0.2$, i.e., $t=5,5.2,5.4,\cdots,9.8,10$. Panel (b) shows the optimum $m$ denoted by $m_{opt}$ over the domain $1\leq m\leq 20$ at which $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest. We have $m_{opt}=1$ for $t\leq 7$ and $m_{opt}=20$ for $t\geq 7.2$. This tells us that if Conjecture 1 is true, then the critical threshold must satisfy $\tau_{c}\leq 7.2$. In Corollary 2 in below, we will present a theoretical upper bound on $\tau_{c}$.
@@
Figure 5: Panel (a) presents plots of $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ in terms of $1\leq m\leq 20$ for $5\leq t\leq 10$ in steps of $0.2$. Panel (b) shows the optimum $m$ denoted by $m_{opt}$ over the domain $1\leq m\leq 20$ at which $\inf_{0<b<1}e^{\Lambda_{m}(t,b)}$ is the smallest. If Conjecture 1 is true, then these plots imply that $\tau_{c}\leq 7.2$.
@@
We continue by stating two corollaries to Theorem 1. For our first corollary, we study the behaviour of the bounds in Theorem 1 as $m$ grows to infinity. This will result in what we refer to as the $m_{\infty}$-inequality.
@@
The first inequality in (52) is trivial. For the second inequality, we note that [^14] If $f_{n}:\mathcal{A}\to\mathbb{R}$ is an arbitrary sequence of functions on a set $\mathcal{A}$ of real numbers, then $\inf_{a\in\mathcal{A}}f_{n}(a)\leq f_{n}(x)$ for every $x\in\mathcal{A}$ and $n\geq 1$. Taking limit inferior from both sides, we get $\liminf_{n\to\infty}\inf_{a\in\mathcal{A}}f_{n}(a)\leq\liminf_{n\to\infty}f_{n}(x)$ for every $x\in\mathcal{A}$. This says $\liminf_{n\to\infty}\inf_{a\in\mathcal{A}}f_{n}(a)$ is a lower bound on the set of numbers $\liminf_{n\to\infty}f_{n}(x)$ for $x\in\mathcal{A}$. Thus, $\liminf_{n\to\infty}\inf_{a\in\mathcal{A}}f_{n}(a)\leq\inf_{x\in\mathcal{A}}\liminf_{n\to\infty}f_{n}(x)$ as the infimum of a set is its largest lower bound.
@@
for all $t\geq 0$. The limit $\lim_{m\to\infty}\Lambda_{m}(t,b)$ exists. To see this, first we derive a different formula for $\theta_{m}(b)$. Recall the Maclaurin series $\ln(1-b)=\sum_{i=1}^{\infty}\frac{b^{i}}{i}$. Then
$\theta_{m}(b)=-\frac{1}{b^{m+1}}\Big(-\sum_{i=m+1}^{\infty}\frac{b^{i}}{i}\Big)=\sum_{i=m+1}^{\infty}\frac{b^{i-m-1}}{i}.$ (54)
Changing the index $i$ to $i+m+1$ results in $\theta_{m}(b)=\sum_{i=0}^{\infty}\frac{b^{i}}{i+m+1}$. We have
@@
where we have used the MacLaurin series for $\ln(1-b)$ one more time. By (57), (58) and (59),
@@
Our second corollary enables us to find an upper bound on the critical threshold $\tau_{c}$ in Conjecture 1.
@@
Let $\hat{\tau}_{c}$ be the unique positive solution for $t$ in the equation
$\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}\exp\Big(-\frac{(\frac{1}{2}-\kappa)t}{\|A\|}\Big)=1,$ (63)
where $\kappa$ is the constant in HWI given in Proposition 1. If Conjecture 1 is true, then $\tau_{c}\leq\hat{\tau}_{c}$.
@@
$g(t)=\liminf_{m\to\infty}\inf_{0<b<1}e^{\Lambda_{m}(t,b)}.$ (65)
By definition, the critical threshold $\tau_{c}$ is given by
$\tau_{c}=\sup\{t>0:g(t)\geq f(t)\}.$ (66)
@@
$\tau_{c}\leq\hat{\tau}_{c}=\sup\Big\{t>0:\Big(1+\frac{t}{n\|A\|}\Big)^{\frac{n}{2}}e^{-\frac{t}{2\|A\|}}\geq e^{-\frac{\kappa t}{\|A\|}}\Big\}.$ (67)
As an application of Corollary 2, let us consider the symmetric matrix $A$ in (51). Using the plots in Fig. 5, we concluded that $\tau_{c}\leq 7.2$. Using Corollary 2, we get the sharper bound $\tau_{c}\leq\hat{\tau}_{c}=6.191$.
II-C Improvement over HWI for both all sufficiently small $t$ and large $t$
Here, we look at the twin inequality to HWI in Theorem 2. For a numerical study, let us consider the symmetric matrix $A$ in (51). Panels (a) and (b) in Fig. 6 compare the bound in (31) with HWI for $0\leq t\leq 2$ and $5\leq t\leq 14$, respectively. We observe that inequality (31) is tighter than HWI for both sufficiently small and large values of the tail parameter $t$. As briefly pointed out in Section I, this advantage can be verified by writing asymptotic expansions for the two functions $\eta_{1}(t)$ and $\eta_{2}(t)$ in (29) and (30), respectively. One can easily verify that
@@ II-C Improvement over HWI for both all sufficiently small $t$ and large $t$
Figure 6: Inequality (31) in Theorem 2 is tighter than HWI for both sufficiently small and large values of $t$. For the symmetric matrix $A$ given in (51), Panel (a) and Panel (b) compare the two bounds for $0\leq t\leq 2$ and $5\leq t\leq 14$, respectively. We have also included the $m_{\infty}$-bound in Corollary 1 for comparison.
@@
This shows that $\min\{\eta_{1}(t),\eta_{2}(t)\}$ behaves similarly to the term $\min\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\}$ for sufficiently small and large values of $t$. Since the coefficient $\kappa^{\prime}\approx 0.187$ that sits behind $\min\{\eta_{1}(t),\eta_{2}(t)\}$ is larger than the coefficient $\kappa\approx 0.145$ behind $\min\{\frac{t^{2}}{\|A\|_{2}^{2}},\frac{t}{\|A\|}\}$ in HWI, it follows that (31) is tighter than HWI for those extreme values of $t$.
@@
Let the matrix $A$ be positive-semidefinite. We present five upper bounds on $\Pr(\boldsymbol{\Delta}>t)$ that depend on $A$ only through the dimension $n$ and the operator norm $\|A\|$. The first candidate is the $m_{\infty}$-bound in Corollary 1 which in fact holds for every symmetric matrix $A$. For the second and third candidates, we consider relaxed versions of HWI and LMI, respectively. Recall the inequality
@@
The exponents in HWI and LMI are increasing in terms of $\|A\|_{2}$. This is clear in case of HWI and its relaxed version in (32) immediately follows. As for LMI in (8), note that $\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2}$ is positive and decreasing in terms of $\|A\|_{2}$ and hence, $(\sqrt{\|A\|_{2}^{2}+2\|A\|t}-\|A\|_{2})^{2}$ becomes increasing in terms of $\|A\|_{2}$. This together with (77) gives
@@
Factoring $\sqrt{n}\|A\|$ in $\sqrt{n\|A\|^{2}+2\|A\|t}-\sqrt{n}\|A\|$ results in the relaxed LMI in (33). As stated in Proposition 4, the $m_{\infty}$-bound is always tighter than both the relaxed HW and the relaxed LM inequalities. One can also consider the possibility of relaxing the augmented and optimal LM inequalities in (47) and (48), respectively. This requires that we replace the ratio $\rho=\frac{t\|A\|}{\|A\|_{2}^{2}}$ in (40) by $\frac{t\|A\|}{(\sqrt{n}\|A\|)^{2}}=\frac{t}{n\|A\|}$. Numerical results indicate that the $m_{\infty}$-bound is still tighter that these relaxed inequalities and hence, we do not pursue this direction.
@@
where $\gamma(a,z)=\int_{0}^{z}v^{a-1}e^{-v}dv$ is the lower incomplete gamma function and $\Gamma(a)=\int_{0}^{\infty}v^{a-1}e^{-v}dv$ is the gamma function. For simplicity and to gain some insight, let us assume that $n$ is an even integer during the rest of this discussion. In this case, $F_{\chi^{2}_{n}}(x)$ admits a closed-form expression given by [^16] One can derive (86) by using the formula $\int v^{k}e^{-v}dv=-\sum_{i=0}^{k}\frac{k!}{i!}v^{i}e^{-v}+C$ and the fact that $\Gamma(k)=(k-1)!$ for positive integer $k$.
@@
Fig. 7 in panels (a) and (b) presents graphs of $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ for $n=8$ and $n=10$, respectively. Approximate values for the numbers $r_{n},r^{\prime}_{n}$ are also provided on these plots. Fig. 8 compares the $m_{\infty}$-bound, the relaxed HW bound, the relaxed LM bound and the $\chi^{2}$-bound for $n=8$ and $n=10$.
Figure 8: Comparison between the $m_{\infty}$-bound, the relaxed HW bound, the relaxed LM bound and the $\chi^{2}$-bound for $n=8$ and $n=10$. The formulas for these four bounds are entirely functions of the dimension $n$ and the ratio $r$ in (88).
@@
$=$ $\frac{1}{\sqrt{e}}\Big(\frac{e(1+r)}{n}\Big)^{\frac{n}{2}}e^{-\frac{r}{2}},\,\,\,\,\,1+r\geq n.$
This is the large deviations bound introduced earlier in (35). It is always weaker than the $m_{\infty}$-bound for every $n\geq 2$. This is shown at the end of the proof for Proposition 4 in Section VIII.